# On continuity of state-dependent utilities

### v1

@@ 3 Statement of the main results
We begin by recalling that we work under the hypotheses of Theorem 3.1. In analogy to the previous section, it is possible to construct a functional $V:{\mathcal{F}}\times\mathcal{L}^{\infty}(\Omega,\mathcal{F})\to\mathbb{R}$ that satisfies (P1), (P2), (P3) and (P4). We also point out that properties (P1-4) are in fact stronger than those required in DM23 , Theorem 3.2777Indeed, using now the properties defined in DM23 , the fact that $V_{\Omega}:\mathcal{L}^{\infty}(\Omega,\mathcal{F})\to\mathbb{R}$ is (${\mathcal{G}}$-Mo), (${\mathcal{G}}$-PC) and (${\mathcal{G}}$-NB) follows trivially from (P3) and (P4), for any sub-$\sigma$-algebra ${\mathcal{G}}$ of ${\mathcal{F}}$. Furthermore, recalling that, by (P1), $V_{A}(0)=0$ we are able to recover also (${\mathcal{G}}$-QL) and (${\mathcal{G}}$-PS). Notice that, for $g_{1},g_{2}\in{\mathcal{L}}^{\infty}(\Omega,{\mathcal{G}}),A\in{\mathcal{G}}$, it is possible to write $V_{\Omega}(g_{1}\mathbf{1}_{A}+g_{2}\mathbf{1}_{A^{c}})=V_{A}(g_{1})+V_{A^{c}}(g_{2})$.. Therefore, $\mathfrak{m}\left(f\middle|{\mathcal{G}}\right)\neq\emptyset$ and for any $g_{1},g_{2}\in\mathfrak{m}\left(f\middle|{\mathcal{G}}\right)$ we have $\{g_{1}\neq g_{2}\}\in{\mathcal{N}}_{\succeq}$. Step 1: We first show the existence of a state dependent utility on $\mathcal{L}^{\infty}(\Omega,\mathcal{G})$ for any $\sigma$-algebra ${\mathcal{G}}\subseteq{\mathcal{F}}$. Consider the functional $V:{\mathcal{F}}\times\mathcal{L}^{\infty}(\Omega,\mathcal{F})\to\mathbb{R}$ defined above. By Proposition 4.1 it follows that $V$ admits a representation of the kind $V_{A}(f)=\int_{A}\mathbb{u}(\omega,f(\omega))\mathrm{d}\mathbb{P}(\omega)$. Let ${\mathcal{G}}\subseteq{\mathcal{F}}$ be an arbitrary sub-$\sigma$-algebra on $\Omega$ and denote by $V^{\mathcal{G}}:{\mathcal{G}}\times\mathcal{L}^{\infty}(\Omega,\mathcal{G})\to\mathbb{R}$ the restriction of the functional $V$ to ${\mathcal{G}}\subseteq{\mathcal{F}}$ and $\mathcal{L}^{\infty}(\Omega,\mathcal{G})$. We stress that, differently from ${\mathcal{F}}$, the $\sigma$-algebra ${\mathcal{G}}$ need not contain at least three disjoint events of positive probability. Analogously to $V$, $V^{\mathcal{G}}$ defines a probability measure $\mathbb{P}_{\mathcal{G}}$ on ${\mathcal{G}}$, which is the restriction of the probability measure $\mathbb{P}$ induced by $V$ to the sub-$\sigma$-algebra ${\mathcal{G}}$. We will hence use only $\mathbb{P}$ in the following, to avoid impractical notations.
@@
Applying Proposition 4.1 to $V^{\mathcal{G}}$ it is possible to construct a ${\mathcal{G}}\otimes\mathbb{B}_{\mathbb{R}}$-measurable function $\mathbb{u}_{\mathcal{G}}:\Omega\times\mathbb{R}\to\mathbb{R}$ such that
@@
It is then immediate to verify by definition of conditional expectation that the random variable $\mathbb{u}_{\mathcal{G}}(\cdot,x)$ is an element of the class $\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(\cdot,x)|{\mathcal{G}}\right]$. To avoid overburdening notation, in the following step 2 we shall write $\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(f)\middle|{\mathcal{G}}\right]$ in place of $\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(\cdot,f(\cdot))|{\mathcal{G}}\right]$ and $\mathbb{u}_{\mathcal{G}}(g)$ instead of $\mathbb{u}_{\mathcal{G}}(\cdot,g(\cdot))$. Moreover, we shall often make the following standard abuse of notation: for any $f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ we write $g=\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(f)|{\mathcal{G}}\right]$ to indicate that $g\in\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(f)|{\mathcal{G}}\right]$ i.e. $g$ is a version of the conditional expectation $\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(f)|{\mathcal{G}}\right]$. Step 2: We show that:
@@
and therefore (15) is verified for ${\mathcal{G}}$-measurable simple functions. We proceed to show it also holds for general functions $g\in\mathcal{L}^{\infty}(\Omega,\mathcal{G})$. Take any sequence of ${\mathcal{G}}$-measurable simple functions $(\underline{g}_{n})_{n}\in\mathcal{L}^{\infty}(\Omega,\mathcal{G})$ such that $\underline{g}_{n}\uparrow_{n}g$ (again, the existence of such functions is guaranteed by standard arguments). Monotonicity of $\mathbb{u}_{\mathcal{G}}(\omega,\cdot)$ ensures that $(\mathbb{u}_{\mathcal{G}}(\omega,\underline{g}_{n}(\omega)))_{n}$ is an increasing sequence for all $\omega\in\Omega$. Finally, pointwise continuity of $\mathbb{u}$ gives
@@
and monotone convergence theorem yields
$\lim_{n}\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(\underline{g}_{n})\middle|{\mathcal{G}}\right]=\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(g)\middle|{\mathcal{G}}\right]$
which provides (15) for general $g\in\mathcal{L}^{\infty}(\Omega,\mathcal{G})$. Step 3: we now study the generalized inverse of the state dependent utility $\mathbb{u}_{\mathcal{G}}$. If on the one hand $(\omega,x)\mapsto\mathbb{u}_{\mathcal{G}}(\omega,x)$ is ${\mathcal{G}}\otimes\mathbb{B}_{\mathbb{R}}$ measurable (by Proposition 4.1), we want the generalized inverse $\Phi_{\mathcal{G}}$ to enjoy the same measurability as $\mathbb{u}_{\mathcal{G}}$. We first extend the domain of $\mathbb{u}_{\mathcal{G}}:\Omega\times\mathbb{R}\to\mathbb{R}$ to $\Omega\times\overline{\mathbb{R}}$ by denoting
@@
Notice that, fixed $\omega\in\Omega$, whenever $x\in\text{Im}_{\omega}(\mathbb{u}_{\mathcal{G}})$ then $\Phi_{\mathcal{G}}(\omega,x)$ is finite. We claim that $\Phi_{\mathcal{G}}(\cdot,x):\Omega\to\overline{\mathbb{R}}$ is ${\mathcal{G}}$-measurable for all $x\in\mathbb{R}$. To see this, let $x\in\mathbb{R}$ be fixed and consider the set $A=(a,+\infty]\in\mathbb{B}_{\overline{\mathbb{R}}}$ with $a>-\infty$. We show that $\left\{\omega\in\Omega:\Phi_{\mathcal{G}}(\omega,x)\in A\right\}\in{\mathcal{G}}$. We observe that, by strict monotonicity of $\mathbb{u}_{\mathcal{G}}$ the latter coincides with $\left\{\omega\in\Omega:x>\mathbb{u}_{\mathcal{G}}(\omega,a)\right\}$ which is an element of ${\mathcal{G}}$ as $\mathbb{u}_{\mathcal{G}}$ is ${\mathcal{G}}$-measurable.
Addittionally, we wish to emphasize that $\Phi_{\mathcal{G}}(\omega,\cdot):\mathbb{R}\to\left[-\infty,+\infty\right]$ is continuous for all $\omega\in\Omega$, and therefore $\Phi_{\mathcal{G}}:\Omega\times\mathbb{R}\to\left[-\infty,+\infty\right]$ is a Carathéodory function (see Aliprantis , Definition 4.50). In particular, we observe that $\mathbb{R}$ with the usual euclidean topology is a separable metrizable space and $\overline{\mathbb{R}}=\left[-\infty,+\infty\right]$ with the standard topology on the extended real numbers is metrizable. Then, by virtue of Aliprantis , Lemma 4.51, $\Phi_{\mathcal{G}}$ is ${\mathcal{G}}\otimes\mathbb{B}_{\mathbb{R}}$ measurable.
Remark 5.1
@@
Taking a version $g$ of $\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(\cdot,f)\middle|{\mathcal{G}}\right]$ we can find $A\in{\mathcal{G}}$ such that $\mathbb{P}(A)=1$ and exploting (15) we may write
@@
Let us define $\tilde{g}=g\mathbf{1}_{A}+0\mathbf{1}_{A^{c}}$ and observe that now $\tilde{g}(\omega)\in\text{Im}_{\omega}(\mathbb{u}_{\mathcal{G}})$ for all $\omega\in\Omega$. Moreover, $\tilde{g}$ is still a version of the conditional expectation of $f$. Recalling that the mapping $\Phi_{\mathcal{G}}(\omega,\cdot):\mathbb{R}\to\overline{\mathbb{R}}$ is strictly increasing on $\text{Im}_{\omega}(\mathbb{u}_{\mathcal{G}})\subseteq\mathbb{R}$ for all $\omega\in\Omega$, and application to (18) yields
@@
Consequently, by (17) $\Phi_{\mathcal{G}}(\omega,\mathbb{u}_{\mathcal{G}}(\omega,x))=x$ for all $\omega\in\Omega$ and all $x\in\mathbb{R}$ and it follows that that $-\left\|f\right\|_{\infty}<\Phi_{\mathcal{G}}(\omega,\tilde{g})<\left\|f\right\|_{\infty}$, for every $\omega\in\Omega$. Hence $\Phi_{\mathcal{G}}\left(\cdot,\tilde{g}\right)\in\mathcal{L}^{\infty}(\Omega,\mathcal{G})$. In particular then we have
@@
We prove that $\Phi_{\mathcal{G}}(\omega,\tilde{g}(\omega))$ is the conditional Chisini mean in that it solves the infinite dimensional system of equations given in (5), for the preference representing functional $V_{\Omega}(f)=\int_{\Omega}\mathbb{u}(\omega,f(\omega))\mathrm{d}\mathbb{P}(\omega)$ defined in Theorem 3.1. Considering an arbitrary set $A\in{\mathcal{G}}\subseteq{\mathcal{F}}$, we have

### v2

@@ 5.3 Proof of Theorem 8
We begin by recalling that we work under the hypotheses of Theorem 7. In analogy to the previous section, it is possible to construct a functional $V:{\mathcal{F}}\times\mathcal{L}^{\infty}(\Omega,\mathcal{F})\to\mathbb{R}$ that satisfies (C1), (C2), (C3) and (C4). We also point out that properties (C1-4) are in fact stronger than those required in [11], Theorem 3.2121212Indeed, using now the properties defined in [11], the fact that $V_{\Omega}:\mathcal{L}^{\infty}(\Omega,\mathcal{F})\to\mathbb{R}$ is (${\mathcal{G}}$-Mo), (${\mathcal{G}}$-PC) and (${\mathcal{G}}$-NB) follows trivially from (C3) and (C4), for any sub-$\sigma$-algebra ${\mathcal{G}}$ of ${\mathcal{F}}$. Furthermore, recalling that, by (C1), $V_{A}(0)=0$ we are able to recover also (${\mathcal{G}}$-QL) and (${\mathcal{G}}$-PS). Notice that, for $g_{1},g_{2}\in{\mathcal{L}}^{\infty}(\Omega,{\mathcal{G}}),A\in{\mathcal{G}}$, it is possible to write $V_{\Omega}(g_{1}\mathbf{1}_{A}+g_{2}\mathbf{1}_{A^{c}})=V_{A}(g_{1})+V_{A^{c}}(g_{2})$.. Therefore, $\mathfrak{m}\left(f\middle|{\mathcal{G}}\right)\neq\emptyset$ and for any $g_{1},g_{2}\in\mathfrak{m}\left(f\middle|{\mathcal{G}}\right)$ we have $\{g_{1}\neq g_{2}\}\in{\mathcal{N}}_{\succeq}$.
Step 1: We first show the existence of a state-dependent utility on $\mathcal{L}^{\infty}(\Omega,\mathcal{G})$ for any $\sigma$-algebra ${\mathcal{G}}\subseteq{\mathcal{F}}$. Consider the functional $V:{\mathcal{F}}\times\mathcal{L}^{\infty}(\Omega,\mathcal{F})\to\mathbb{R}$ defined above. By Proposition 11 it follows that $V$ admits a representation of the kind $V_{A}(f)=\int_{A}\mathbb{u}(\omega,f(\omega))\mathrm{d}\mathbb{P}(\omega)$. Let ${\mathcal{G}}\subseteq{\mathcal{F}}$ be an arbitrary sub-$\sigma$-algebra on $\Omega$ and denote by $V^{\mathcal{G}}:{\mathcal{G}}\times\mathcal{L}^{\infty}(\Omega,\mathcal{G})\to\mathbb{R}$ the restriction of the functional $V$ to ${\mathcal{G}}\subseteq{\mathcal{F}}$ and $\mathcal{L}^{\infty}(\Omega,\mathcal{G})$. We stress that, differently from ${\mathcal{F}}$, the $\sigma$-algebra ${\mathcal{G}}$ need not contain at least three disjoint events of positive probability. Analogously to $V$, $V^{\mathcal{G}}$ defines a probability measure $\mathbb{P}_{\mathcal{G}}$ on ${\mathcal{G}}$, which is the restriction of the probability measure $\mathbb{P}$ induced by $V$ to the sub-$\sigma$-algebra ${\mathcal{G}}$. We will hence use only $\mathbb{P}$ in the following, to avoid impractical notations.
@@
Applying Proposition 11 to $V^{\mathcal{G}}$ it is possible to construct a ${\mathcal{G}}\otimes\mathbb{B}_{\mathbb{R}}$-measurable function $\mathbb{u}_{\mathcal{G}}:\Omega\times\mathbb{R}\to\mathbb{R}$ such that
@@
It is then immediate to verify by definition of conditional expectation that the random variable $\mathbb{u}_{\mathcal{G}}(\cdot,x)$ is an element of the class $\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(\cdot,x)|{\mathcal{G}}\right]$.
To avoid overburdening notation, in the following step 2 we shall write $\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(f)\middle|{\mathcal{G}}\right]$ in place of $\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(\cdot,f(\cdot))|{\mathcal{G}}\right]$ and $\mathbb{u}_{\mathcal{G}}(g)$ instead of $\mathbb{u}_{\mathcal{G}}(\cdot,g(\cdot))$. Moreover, we shall often make the following standard abuse of notation: for any $f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ we write $g=\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(f)|{\mathcal{G}}\right]$ to indicate that $g\in\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(f)|{\mathcal{G}}\right]$ i.e. $g$ is a version of the conditional expectation $\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(f)|{\mathcal{G}}\right]$.
Step 2: We show that:
@@
and therefore (22) is verified for ${\mathcal{G}}$-measurable simple functions. We proceed by showing that the property also holds for general functions $g\in\mathcal{L}^{\infty}(\Omega,\mathcal{G})$. Take any sequence of ${\mathcal{G}}$-measurable simple functions $(\underline{g}_{n})_{n}\in\mathcal{L}^{\infty}(\Omega,\mathcal{G})$ such that $\underline{g}_{n}\uparrow_{n}g$ (again, the existence of such functions is guaranteed by standard arguments). Monotonicity of $\mathbb{u}_{\mathcal{G}}(\omega,\cdot)$ ensures that $(\mathbb{u}_{\mathcal{G}}(\omega,\underline{g}_{n}(\omega)))_{n}$ is an increasing sequence for all $\omega\in\Omega$. Finally, pointwise continuity of $\mathbb{u}$ gives
@@
and monotone convergence theorem yields (22) for general $g\in\mathcal{L}^{\infty}(\Omega,\mathcal{G})$ as
$\lim_{n}\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(\underline{g}_{n})\middle|{\mathcal{G}}\right]=\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(g)\middle|{\mathcal{G}}\right].$
Step 3: we now study the generalized inverse of the state-dependent utility $\mathbb{u}_{\mathcal{G}}$. If on the one hand $(\omega,x)\mapsto\mathbb{u}_{\mathcal{G}}(\omega,x)$ is ${\mathcal{G}}\otimes\mathbb{B}_{\mathbb{R}}$ measurable (by Proposition 11), on the other we would like the generalized inverse $\Phi_{\mathcal{G}}$ to enjoy the same measurability as $\mathbb{u}_{\mathcal{G}}$. We first extend the domain of $\mathbb{u}_{\mathcal{G}}:\Omega\times\mathbb{R}\to\mathbb{R}$ to $\Omega\times\overline{\mathbb{R}}$ by denoting
@@
Notice that, fixed $\omega\in\Omega$, whenever $x\in\text{Im}_{\omega}(\mathbb{u}_{\mathcal{G}})$ then $\Phi_{\mathcal{G}}(\omega,x)$ is finite.
We claim that $\Phi_{\mathcal{G}}(\cdot,x):\Omega\to\overline{\mathbb{R}}$ is ${\mathcal{G}}$-measurable for all $x\in\mathbb{R}$. To see this, let $x\in\mathbb{R}$ be fixed and consider the set $A=(a,+\infty]\in\mathbb{B}_{\overline{\mathbb{R}}}$ with $a>-\infty$. We show that $\left\{\omega\in\Omega:\Phi_{\mathcal{G}}(\omega,x)\in A\right\}\in{\mathcal{G}}$. We observe that, by strict monotonicity of $\mathbb{u}_{\mathcal{G}}$ the latter coincides with $\left\{\omega\in\Omega:x>\mathbb{u}_{\mathcal{G}}(\omega,a)\right\}$ which is an element of ${\mathcal{G}}$ as $\mathbb{u}_{\mathcal{G}}$ is ${\mathcal{G}}$-measurable.
Additionally, we wish to emphasize that $\Phi_{\mathcal{G}}(\omega,\cdot):\mathbb{R}\to\left[-\infty,+\infty\right]$ is continuous for all $\omega\in\Omega$, and therefore $\Phi_{\mathcal{G}}:\Omega\times\mathbb{R}\to\left[-\infty,+\infty\right]$ is a Carathéodory function (see [1], Definition 4.50). In particular, we observe that $\mathbb{R}$ with the usual euclidean topology is a separable metrizable space and $\overline{\mathbb{R}}=\left[-\infty,+\infty\right]$ with the standard topology on the extended real numbers is metrizable. Then, by virtue of [1], Lemma 4.51, $\Phi_{\mathcal{G}}$ is ${\mathcal{G}}\otimes\mathbb{B}_{\mathbb{R}}$ measurable.
Remark 16.
@@
Taking a version $g$ of $\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(\cdot,f)\middle|{\mathcal{G}}\right]$ we can find $A\in{\mathcal{G}}$ such that $\mathbb{P}(A)=1$ and exploiting (22) we may write
@@
Let us define $\tilde{g}=g\mathbf{1}_{A}+0\mathbf{1}_{A^{c}}$ and observe that now $\tilde{g}(\omega)\in\text{Im}_{\omega}(\mathbb{u}_{\mathcal{G}})$ for all $\omega\in\Omega$. Moreover, $\tilde{g}$ is still a version of the conditional expectation of $f$. Recalling that the mapping $\Phi_{\mathcal{G}}(\omega,\cdot):\mathbb{R}\to\overline{\mathbb{R}}$ is strictly increasing on $\text{Im}_{\omega}(\mathbb{u}_{\mathcal{G}})\subseteq\mathbb{R}$ for all $\omega\in\Omega$, an application to (25) yields
@@
Consequently, by (24) $\Phi_{\mathcal{G}}(\omega,\mathbb{u}_{\mathcal{G}}(\omega,x))=x$ for all $\omega\in\Omega$ and all $x\in\mathbb{R}$ and it follows that that $-\left\|f\right\|_{\infty}<\Phi_{\mathcal{G}}(\omega,\tilde{g})<\left\|f\right\|_{\infty}$, for every $\omega\in\Omega$. Hence $\Phi_{\mathcal{G}}\left(\cdot,\tilde{g}\right)\in\mathcal{L}^{\infty}(\Omega,\mathcal{G})$. In particular we have
@@
We prove that $\Phi_{\mathcal{G}}(\omega,\tilde{g}(\omega))$ is the conditional Chisini mean in that it solves the infinite dimensional system of equations given in (10), for the preference representing functional $V_{\Omega}(f)=\int_{\Omega}\mathbb{u}(\omega,f(\omega))\mathrm{d}\mathbb{P}(\omega)$ defined in Theorem 7. Considering an arbitrary set $A\in{\mathcal{G}}\subseteq{\mathcal{F}}$, we have